# Tap into Reality: Understanding the Impact of Interactions on Presence and Reaction Time in Mixed Reality

### v1

@@ 7 Conclusion
In conclusion, this study provides a comprehensive analysis of how interaction and task types influence presence and reaction time, offering valuable insights for enhancing user experience in MR environments. The confirmation of our hypotheses strengthens the argument that direct interactions, which are inherently more immersive, enhance presence compared to symbolic interactions, reflecting the intuitive nature of direct manipulation in VE. The study confirms that changes in interaction type alone can meaningfully alter presence and reaction time. This finding has practical implications for MR developers, highlighting the importance of careful interaction designs to enhance user experience. In summary, these findings expand our understanding of MR interactions and point to reaction time as a promising metric for assessing presence in future MR applications, warranting further exploration.

### v2

@@ 7 Conclusion
In conclusion, this study provides a comprehensive analysis of how interaction and task types influence presence and reaction time, offering valuable insights for enhancing user experience in MR environments. The confirmation of H1 and H2 highlights the immersive potential of direct interactions for tasks requiring high immersion compared to symbolic interactions. Future work should examine additional symbolic interactions to provide broader generalizability. The study confirms that changes in interaction type alone can meaningfully alter presence and reaction time. This finding has practical implications for MR developers, highlighting the importance of careful interaction designs to enhance user experience. In summary, these findings expand our understanding of MR interactions and point to reaction time as a promising metric for assessing presence in future MR applications, warranting further exploration.
Acknowledgements.
We thank the anonymous IEEE VR 2025 reviewers for their insightful comments and feedback. The research reported in this paper was sponsored by the National Science Foundation under award 2237485.